Doris invested some money at 7% and some money at 5% She invested 6,000 more at 5% than she did at 7% Her total yearly interest from the two investments was 600 How much did Doris invest at each rate?

Found 2 solutions by jorel1380, checkley79:
Answer by jorel1380(2518)

.07x+.05(x+6000)=600
7x+5(x+6000)=60000
7x+5x+30000=60000
12x=30000
x=30000/12=2500
x+6000=8500..

Answer by checkley79(3056)

.07x+.05(x+6,000)=600
.07x+.05x+300=600
.12x=600-300
.12x=300
x=300/.12
x=$2,500 amount invested @ 7%.
2,500+6,000=$ 8,500 amount invested @ 5%.
Proof:
.07*2,500+.05*8,500=600
175+425=600
600=600